# A convergence proof for a finite element discretization of Chorin’s projection method of the incompressible Navier-Stokes equations

### v1

@@ 4. Convergence analysis of the scheme
$\begin{split}\left|\int_{0}^{T}(viii)dt\right|&=\left|\int_{0}^{T}(p_{h},\operatorname{div}(\mathcal{P}_{\mathbb{U}_{h}}v-v))dt\right|\\ &\leq\left\lVert p_{h}\right\rVert_{L^{2}([0,T]\times\Omega)}\left\lVert\operatorname{div}(\mathcal{P}_{\mathbb{U}_{h}}v-v)\right\rVert_{L^{2}([0,T]\times\Omega)}\\ &\leq\frac{1}{\sqrt{\Delta t}}\left(\sum_{m=0}^{N-1}\left\lVert\widetilde{u}^{m+1}_{h}-u^{m+1}_{h}\right\rVert_{L^{2}}^{2}\right)^{\frac{1}{2}}\left(\int_{0}^{T}|v-\mathcal{P}_{\mathbb{U}_{h}}v|_{H^{1}}^{2}dt\right)^{\frac{1}{2}}\\ &\leq\frac{C(E_{h}^{0}+\left\lVert f\right\rVert_{L^{2}([0,T]\times\Omega)}^{2})^{1/2}h}{\sqrt{\Delta t}}\left\lVert v\right\rVert_{L^{2}(0,T;H^{2}(\Omega))},\end{split}$
similar to the estimate (4.11). Clearly, also this term goes to zero if $h=o(\sqrt{\Delta t})$. Finally, for the ninth term, we have, since $f_{h}=\mathcal{P}_{\mathbb{U}_{h}}f^{m}$,
@@
(4.22) $\int_{0}^{T}\left[(u,\partial_{t}v)+b(u,v,u)-\mu(\nabla u,\nabla v)+(f,v)\right]dt+(u_{0},v(0,\cdot))=0,$
for any divergence free $v\in C^{\infty}_{c}([0,T)\times\Omega))$. Passing to the limit in the discrete energy balance (3.18) and using the properties of weak convergence, we obtain that the limit $u$ satisfies for almost any $t\in[0,T]$,
@@ 5. Convergence of projection method with projection step as a Poisson problem
While the discretization of the projection step in (3.2) could be considered as a Darcy problem, one can consider a discretization as a Poisson problem. In this case, one can obtain stability and convergence of a fully discrete scheme for the incremental version of the projection method. This was already done in [12] by the use of a special interpolation operator which was constructed explicitly for the lowest degree Taylor-Hood finite elements. No inf-sup condition such as (3.8) needed to be assumed for this to work. We will show here that the proof can be modified to prove that the linearly implicit discretization of the incremental projection method below converges without the need for this interpolation operator and therefore independently of the degree of the finite elements. This requires an adaption of the compactness arguments in [12] but also shortens the proof significantly. We will outline the compactness arguments here and sketch the rest of the proof as it can be achieved in a similar way to the proof of convergence of Chorin’s version of the projection method in the previous sections or the proof in [12]. We consider the following discretization of (1.1): We use the same notation and trilinear form $b$ as in Section 3.2 with the following modifications. We will seek the intermediate velocity $\widetilde{u}_{h}^{m+1}$ in a space $\mathbb{U}_{h}\subset H^{1}_{0}(\Omega)$, the pressure $p_{h}^{m+1}$ in a space $\mathbb{P}_{h}\subset H^{1}(\Omega)\cap L^{2}_{0}(\Omega)$ and the final velocity $u_{h}^{m+1}$ in the space $\mathbb{Y}_{h}=\mathbb{U}_{h}+\nabla\mathbb{P}_{h}$. Note that now $\mathbb{P}_{h}\subset H^{1}(\Omega)$ and therefore $\nabla\mathbb{P}_{h}$ is well-defined (though possibly containing discontinuous functions). Also note that the final velocity $u_{h}^{m+1}$ may not satisfy homogeneous Dirichlet boundary conditions in this case. We require both the $L^{2}$-orthogonal projections $\mathcal{P}_{\mathbb{U}_{h}}:L^{2}(\Omega)\to\mathbb{U}_{h}$ and $\mathcal{P}_{\mathbb{P}_{h}}:L^{2}(\Omega)\to\mathbb{P}_{h}$ defined by (3.6) to satisfy the properties (3.7). In contrast to the previous sections, $\mathbb{P}_{h}$ is now a subspace of $H^{1}(\Omega)$ and the finite elements therefore may have more regularity. As in [12], we define the space
$(\mathcal{P}_{\mathbb{V}_{h}}v,w)=(v,w),\quad\forall w\in\mathbb{V}_{h}.$
thus
Lemma 5.1.
We have that
Proof.
implying the claim. ∎
@@ 1. Step 1
    (Prediction step): For any $m\geq 0$, given $u_{h}^{m}\in\mathbb{Y}_{h}$, $p_{h}^{m}\in\mathbb{P}_{h}$, find $\widetilde{u}_{h}^{m+1}\in\mathbb{U}_{h}$, such that for all $v\in\mathbb{U}_{h}$
@@ 2. Step 2
The solvability of the scheme (5.3)–(5.4) is shown in the same way as for the nonincremental Darcy version of the scheme above in Lemma 3.4, or see also [12, Lemma 2.1]. Furthermore, the scheme satisfies the discrete energy estimate
@@
As in Remark 3.6, this results in a uniform in $h,\Delta t>0$ bound on the energy after applying the discrete Grönwall inequality, Lemma A.1. As in the previous section, we denote piecewise constant and piecewise linear interpolations of the approximations $u_{h}^{m}$, $\widetilde{u}_{h}^{m}$ and $p_{h}^{m}$ by (4.1)–(4.5).
@@ 5.1. Convergence of the numerical scheme
For this scheme, we do not have $\operatorname{div}u_{h}=0$ almost everywhere, we only have the weak form $(u^{m+1}_{h},\nabla q)=0$ for all $q\in\mathbb{P}_{h}$. Since $u^{m+1}_{h}$ is now sought in the space $\mathbb{Y}_{h}=\mathbb{X}_{h}+\nabla\mathbb{P}_{h}$ instead of in $\mathbb{U}_{h}$, we also do not know if $u^{m+1}_{h}$ is zero on the boundary of the domain and therefore we cannot integrate by parts in the weak form to conclude that the divergence is zero almost everywhere. In order to deal with this issue, we need to derive an alternative version of Lemma 4.3 by Lions which is similar to Lemma 3.5 in [12], except that different spaces are used and no interpolation operator $\Pi_{N}$ is needed besides the standard $L^{2}$-orthogonal projections (3.6), and the orthogonal projection $\mathcal{P}_{\mathbb{V}_{h}}$.
First, from the energy balance, Lemma 5.4, we obtain the following uniform a priori estimates for the sequences $\{\bar{u}_{h}\}_{h>0}$ and $\{\widetilde{u}_{h}\}_{h>0}$:
$\{\bar{u}_{h}\}_{h>0}\subset L^{\infty}(0,T;L^{2}(\Omega)),$
@@
(5.6) $\bar{u}_{h}\overset{\star}{\rightharpoonup}\bar{u},$ $\quad\text{in }\,L^{\infty}(0,T;L^{2}(\Omega)),$
(5.7) $\widetilde{u}_{h}\overset{\star}{\rightharpoonup}\widetilde{u}$, $\quad\text{in }\,L^{\infty}(0,T;L^{2}(\Omega)),$
@@
Due to the discrete energy estimate (5.5), Lemma 4.2 still applies to the sequences $\{\bar{u}_{h}\}_{h>0}$ and $\{\widetilde{u}_{h}\}_{h>0}$ and they have the same weak limit $u\in L^{2}(0,T;H^{1}_{0}(\Omega))\cap L^{\infty}(0,T;L^{2}(\Omega))$. The limit is weakly divergence free:
@@
Any weak limit $u$ of the sequence $\{\bar{u}_{h}\}_{h>0}$ satisfies
$(u,\nabla q)=0,\quad\forall\,q\in H^{1}(\Omega).$
Proof.
Since $u$ is the weak limit of $\{\bar{u}_{h}\}_{h>0}$ , we have
$(u,\nabla q)=\lim_{h\to 0}(\bar{u}_{h},\nabla q),$
for any $q\in H^{1}(\Omega)$ for a.e. $t\in[0,T]$. Then we can write
$|(\bar{u}_{h},\nabla q)|=|(\bar{u}_{h},\nabla(q-\mathcal{P}_{\mathbb{P}_{h}}q))|\leq\left\lVert\bar{u}_{h}\right\rVert_{L^{2}}\left\lVert\nabla(q-\mathcal{P}_{\mathbb{P}_{h}}q)\right\rVert_{L^{2}}\stackrel{{\scriptstyle h\to 0}}{{\longrightarrow}}0,$
since $\bar{u}_{h}\in L^{\infty}([0,T];L^{2}(\Omega))$ and $\mathcal{P}_{\mathbb{P}_{h}}q\to q$ in $H^{1}(\Omega)$ as $h\to 0$ for any $q\in H^{1}(\Omega)$. Thus, in the limit $(u,\nabla q)=0$ for any $q\in H^{1}(\Omega)$. ∎
In order to pass to the limit in the weak formulation of the scheme, we need to derive strong convergence of the approximations of the velocity field in $L^{2}([0,T]\times\Omega)$. To do so, we need to prove an estimate on the continuity in time of $\widetilde{u}_{h}$ with values in a weak space (this lemma is a slight adaption to a different space of Lemma 3.3 in [12]):
Lemma 5.7.
Denote $Z=(H^{1}_{\operatorname{div}}(\Omega)\cap H^{2}(\Omega))^{*}$. Assume that $h\leq C\Delta t$. Then the aproximations $\widetilde{u}_{h}$ computed by (5.3)–(5.4) satisfy for any $0\leq\tau<T$,
@@
Given $\tau>0$ and $t\in[0,T-\tau]$, let $m_{1}\in\mathbb{N}$ such that $\widetilde{u}_{h}(t)=u_{h}^{m_{1}}$ and $m_{2}\in\mathbb{N}$ such that $\widetilde{u}_{h}(t+\tau)=\widetilde{u}_{h}^{m_{2}+1}$. Then
@@
We take a test function $v\in H^{1}_{\operatorname{div}}(\Omega)\cap H^{2}(\Omega)$, and plug in the scheme (5.3)–(5.4),
@@
$\leq\left\lVert\widetilde{u}^{m}_{h}\right\rVert_{L^{2}}\left\lVert\nabla\widetilde{u}_{h}^{m+1}\right\rVert_{L^{2}}\left\lVert\mathcal{P}_{\mathbb{U}_{h}}v\right\rVert_{L^{\infty}}$
$\leq\left\lVert\widetilde{u}^{m}_{h}\right\rVert_{L^{2}}\left\lVert\nabla\widetilde{u}_{h}^{m+1}\right\rVert_{L^{2}}\left\lVert v\right\rVert_{L^{\infty}}$
$\leq\left\lVert\widetilde{u}^{m}_{h}\right\rVert_{L^{2}}\left\lVert\nabla\widetilde{u}_{h}^{m+1}\right\rVert_{L^{2}}\left\lVert v\right\rVert_{H^{2}},$
@@
using (3.7b). For the third term, III, we have, using that $v$ is divergence free, the properties of the $L^{2}$-projection, (3.7), and that by the energy estimate $\left\lVert\nabla p\right\rVert_{L^{\infty}(0,T;L^{2}(\Omega))}\leq C\Delta t^{-2}$,
@@
$\leq C(\left\lVert\nabla p^{m}_{h}\right\rVert_{L^{2}(\Omega)}+\left\lVert\nabla p^{m-1}_{h}\right\rVert_{L^{2}(\Omega)})h^{2}\left\lVert v\right\rVert_{H^{2}(\Omega)}$
$\leq C\frac{h^{2}}{\Delta t^{2}}\left\lVert v\right\rVert_{H^{2}(\Omega)}$
$\leq C\left\lVert v\right\rVert_{H^{2}(\Omega)},$
under the condition that $h\leq C\Delta t$. Finally, for the last term, IV, we have
@@
Combining the estimates I – IV, we have, after taking the supremum over $v\in H^{1}_{\operatorname{div}}(\Omega)\cap H^{2}(\Omega)$,
$\left\lVert\widetilde{u}_{h}(t+\tau)-\widetilde{u}_{h}(t)\right\rVert_{Z}\leq C\Delta t\sum_{m=m_{1}}^{m_{2}}\left(\left\lVert\widetilde{u}^{m}_{h}\right\rVert_{L^{2}}\left\lVert\nabla\widetilde{u}_{h}^{m+1}\right\rVert_{L^{2}}+\left\lVert\nabla\widetilde{u}^{m+1}_{h}\right\rVert_{L^{2}(\Omega)}+1+\left\lVert f^{m}\right\rVert_{L^{2}(\Omega)}\right).$
@@
$\begin{split}&\int_{0}^{T-\tau}\left\lVert\widetilde{u}_{h}(t+\tau)-\widetilde{u}_{h}(t)\right\rVert^{2}_{Z}dt\\ &\leq C(\tau+\Delta t)\int_{0}^{T-\tau}\int_{t}^{t+\tau}\left(\left\lVert\widetilde{u}_{h}(s-\Delta t)\right\rVert_{L^{2}}^{2}\left\lVert\nabla\widetilde{u}_{h}(s)\right\rVert^{2}_{L^{2}}+\left\lVert\nabla\widetilde{u}_{h}\right\rVert^{2}_{L^{2}(\Omega)}+1+\left\lVert f_{h}\right\rVert^{2}_{L^{2}(\Omega)}\right)dsdt\\ &\leq C(\tau+\Delta t)\tau\int_{0}^{T-\tau}\left(\left\lVert\widetilde{u}_{h}(s-\Delta t)\right\rVert_{L^{2}}^{2}\left\lVert\nabla\widetilde{u}_{h}(s)\right\rVert^{2}_{L^{2}}+\left\lVert\nabla\widetilde{u}_{h}\right\rVert^{2}_{L^{2}(\Omega)}+1+\left\lVert f_{h}\right\rVert^{2}_{L^{2}(\Omega)}\right)ds\\ &\leq C(\tau+\Delta t)\tau.\end{split}$
Finally, we need the following Lions-like lemma, which will then allow us to prove a stronger time continuity estimate and apply Simon’s compactness criterion:
Lemma 5.8.
Denote $Z=(H^{2}(\Omega)\cap H^{1}_{\operatorname{div}}(\Omega))^{*}$. Then for any $\eta>0$ there exists a constant $C_{\eta}>0$ and $h_{\eta}>0$ such that for all $h\leq h_{\eta}$ we have
(5.10) $\left\lVert\mathcal{P}_{\mathbb{V}_{h}}w\right\rVert_{L^{2}(\Omega)}\leq\eta\left\lVert w\right\rVert_{H^{1}_{0}(\Omega)}+C_{\eta}\left\lVert w\right\rVert_{Z}\quad\forall w\in H^{1}_{0}(\Omega).$
Proof.
$\begin{split}\left\lVert w\right\rVert_{Z}&=\sup_{\varphi\in H^{2}\cap H^{1}_{\operatorname{div}}(\Omega)}\frac{|(w,\varphi)|}{\left\lVert\varphi\right\rVert_{H^{2}\cap H_{\operatorname{div}}^{1}}}\\ &=\sup_{\varphi\in H^{2}\cap H^{1}_{0}(\Omega)}\frac{|(w,\mathcal{P}\varphi)|}{\left\lVert\mathcal{P}\varphi\right\rVert_{H^{2}\cap H_{0}^{1}}}\\ &=\sup_{\varphi\in H^{2}\cap H^{1}_{0}(\Omega)}\frac{|(\mathcal{P}w,\varphi)|}{\left\lVert\mathcal{P}\varphi\right\rVert_{H^{2}\cap H^{1}_{0}}}\\ &\geq\sup_{\varphi\in H^{2}\cap H^{1}_{0}(\Omega)}\frac{|(\mathcal{P}w,\varphi)|}{\left\lVert\varphi\right\rVert_{H^{2}\cap H^{1}_{0}}}\\ &=\left\lVert\mathcal{P}w\right\rVert_{(H^{2}\cap H^{1}_{0})^{*}}\end{split}$
$\left\lVert\mathcal{P}w\right\rVert_{(H^{2}\cap H^{1}_{0})^{*}}=0,$
i.e., $\mathcal{P}w=0$ in $(H^{2}\cap H^{1}_{0})^{*}$. That means that
$(\mathcal{P}w,\varphi)=0,\quad\forall\,\varphi\in H^{2}(\Omega)\cap H^{1}_{0}(\Omega).$
Since $\mathcal{P}w\in L^{2}(\Omega)$, we can extend this identity to $\varphi\in L^{2}(\Omega)$ and conclude that $\left\lVert\mathcal{P}w\right\rVert_{L^{2}(\Omega)}=0$. Next, we want to show that $\mathcal{P}_{N}w_{N}\to\mathcal{P}w$ in $L^{2}(\Omega)$ which would then imply a contraction, since $\left\lVert\mathcal{P}_{N}w_{N}\right\rVert_{L^{2}}=1\neq 0=\left\lVert\mathcal{P}w\right\rVert_{L^{2}}$. We start by showing that $\mathcal{P}_{N}w_{N}\rightharpoonup\mathcal{P}w$ in $L^{2}(\Omega)$. Let us take $\varphi\in L^{2}(\Omega)$ and decompose using the Helmholtz-decomposition as $\varphi=\mathcal{P}\varphi+\nabla g$ where $g\in H^{1}(\Omega)$. Then we have, using Lemma 5.1 and that $w_{N}\to w$ in $L^{2}(\Omega)$,
Combining these two insights, we obtain
which yields the desired contradiction. ∎
Next, we use this result to show uniform time continuity of $\widetilde{u}_{h}$ in $L^{2}([0,T]\times\Omega)$ which we can then use together with Simon’s version of the Aubin-Lions-Simon lemma [27, Theorem 1], stated in Theorem B.1, to conclude precompactness of the sequence $\{\widetilde{u}_{h}\}_{h>0}$ in $L^{2}([0,T]\times\Omega)$. The following lemma and its proof closely follow [12, Lemma 3.6]. We include it here both for completeness and to illustrate that it can be established using Lemma 5.8, instead of the version employed in [12].

### v2

@@ 4. Convergence analysis of the scheme
(4.28) $\begin{split}\left|\int_{0}^{T}(viii)dt\right|&=\left|\int_{0}^{T}(p_{h},\operatorname{div}(\mathcal{P}_{\mathbb{U}_{h}}v-v))dt\right|\\ &\leq\left\lVert p_{h}\right\rVert_{L^{2}([0,T]\times\Omega)}\left\lVert\operatorname{div}(\mathcal{P}_{\mathbb{U}_{h}}v-v)\right\rVert_{L^{2}([0,T]\times\Omega)}\\ &\leq\frac{1}{\sqrt{\Delta t}}\left(\sum_{m=0}^{N-1}\left\lVert\widetilde{u}^{m+1}_{h}-u^{m+1}_{h}\right\rVert_{L^{2}}^{2}\right)^{\frac{1}{2}}\left(\int_{0}^{T}|v-\mathcal{P}_{\mathbb{U}_{h}}v|_{H^{1}}^{2}dt\right)^{\frac{1}{2}}\\ &\leq\frac{C(E_{h}^{0}+\left\lVert f\right\rVert_{L^{2}([0,T]\times\Omega)}^{2})^{1/2}h^{k}}{\sqrt{\Delta t}}\left\lVert v\right\rVert_{L^{2}(0,T;H^{k+1}(\Omega))},\end{split}$
similar to the estimate (4.10), and where we have used (3.7d). Clearly, also this term goes to zero if $h^{k}=o(\sqrt{\Delta t})$. Finally, for the ninth term, we have, since $f_{h}=\mathcal{P}_{\mathbb{U}_{h}}f^{m}$,
@@
(4.29) $\int_{0}^{T}\left[-(u,\partial_{t}v)+b(u,u,v)+\mu(\nabla u,\nabla v)-(f,v)\right]dt=0,$
for any divergence free $v\in C^{\infty}_{c}((0,T)\times\Omega))$. Passing to the limit in the discrete energy balance (3.17) and using the properties of weak convergence, we obtain that the limit $u$ satisfies for almost any $t\in[0,T]$,
@@ 5. Convergence of projection method with projection step as a Poisson problem
While the discretization of the projection step in (3.2) could be considered as a Darcy problem, one can also consider a discretization as a Poisson problem. In this case, one can obtain stability and convergence of a fully discrete scheme for the incremental version of the projection method. This was already done in [12] by the use of a special interpolation operator which was constructed explicitly for the lowest degree Taylor-Hood finite elements. No inf-sup condition such as (3.8) needed to be assumed for this to work. We will show here that the proof can be modified to prove that the linearly implicit discretization of the incremental projection method below converges without the need for this interpolation operator and therefore independently of the degree of the finite elements. This requires an adaption of the compactness arguments in [12] but also shortens the proof significantly. We will outline the compactness arguments here and sketch the rest of the proof as it can be achieved in a similar way to the proof of convergence of Chorin’s version of the projection method in the previous sections or alternatively the proof in [12]. We consider the following discretization of (1.1): We use the same notation and trilinear form $b$ as in Section 3.2 with the following modifications. We will seek the intermediate velocity $\widetilde{u}_{h}^{m+1}$ in a space $\mathbb{U}_{h}\subset H^{1}_{0}(\Omega)$, the pressure $p_{h}^{m+1}$ in a space $\mathbb{P}_{h}\subset H^{1}(\Omega)\cap L^{2}_{0}(\Omega)$ and the final velocity $u_{h}^{m+1}$ in the space $\mathbb{Y}_{h}=\mathbb{U}_{h}+\nabla\mathbb{P}_{h}$. Note that now $\mathbb{P}_{h}\subset H^{1}(\Omega)$ and therefore $\nabla\mathbb{P}_{h}$ is well-defined (though possibly containing discontinuous functions). Also note that the final velocity $u_{h}^{m+1}$ may not satisfy homogeneous Dirichlet boundary conditions in this case. We again require both the $L^{2}$-orthogonal projections $\mathcal{P}_{\mathbb{U}_{h}}:L^{2}(\Omega)\to\mathbb{U}_{h}$ and $\mathcal{P}_{\mathbb{P}_{h}}:L^{2}(\Omega)\to\mathbb{P}_{h}$ defined by (3.6) to satisfy the properties (3.7) for possibly different $k$.
@@ 1. Step 1
    (Prediction step): For any $m\geq 0$, given $\widetilde{u}_{h}^{m}\in\mathbb{U}_{h}$, $u_{h}^{m}\in\mathbb{Y}_{h}$, $p_{h}^{m}\in\mathbb{P}_{h}$, find $\widetilde{u}_{h}^{m+1}\in\mathbb{U}_{h}$, such that for all $v\in\mathbb{U}_{h}$,
@@ 2. Step 2
The solvability of the scheme (5.1)–(5.2) is shown in the same way as for the nonincremental Darcy version of the scheme above in Lemma 3.3, alternatively see also [12, Lemma 2.1]. Furthermore, the scheme satisfies the discrete energy estimate
@@
As in Remark 3.5, this results in a uniform in $h,\Delta t>0$ bound on the energy after applying the discrete Grönwall inequality, Lemma A.1. Next, we define the piecewise constant interpolations in time of the approximations $u_{h}^{m}$, $\widetilde{u}_{h}^{m}$ and $p_{h}^{m}$ by
(5.4) $u_{h}(t)$ $=u_{h}^{m},\qquad$ $t\in[t^{m},t^{m+1}),$
(5.6) $p_{h}(t)$ $=p_{h}^{m+1},\qquad$ $t\in(t^{m},t^{m+1}],$
for $m=0,1,2,\dots$ with
$\widetilde{u}_{h}(r)=u^{0}_{h},\quad\text{and}\quad p_{h}(r)=p_{h}^{0},\quad\text{for }\,r\leq 0.$
@@ 5.1. Convergence of the numerical scheme
For the proof of convergence of this scheme, we enounter similar issues as for the previous scheme that was based on a Darcy formulation of the projection step. On the one hand, we have uniform bounds on $\widetilde{u}_{h}$ in $L^{2}(0,T;H^{1}_{0}(\Omega))$ on the other hand, $u_{h}$ is approximately weakly divergence free, but we cannot say the same for $\widetilde{u}_{h}$ directly. Furthermore, we have very weak bounds on the pressure approximation.
First, from the energy balance, Lemma 5.3, we obtain the following uniform a priori estimates for the sequences $\{{u}_{h}\}_{h>0}$ and $\{\widetilde{u}_{h}\}_{h>0}$:
$\{{u}_{h}\}_{h>0}\subset L^{\infty}(0,T;L^{2}(\Omega)),$
@@
(5.8) ${u}_{h}\overset{\star}{\rightharpoonup}{u},\quad\widetilde{u}_{h}\overset{\star}{\rightharpoonup}\widetilde{u}$, $\quad\text{in }\,L^{\infty}(0,T;L^{2}(\Omega)),$
@@
Due to the discrete energy estimate (5.3), Lemma 4.2 still applies to the sequences $\{{u}_{h}\}_{h>0}$ and $\{\widetilde{u}_{h}\}_{h>0}$ and they have the same weak limit $u\in L^{2}(0,T;H^{1}_{0}(\Omega))\cap L^{\infty}(0,T;L^{2}(\Omega))$. It also follows just as in Lemma 4.3 that the limit is weakly divergence free. However, in this case, we cannot conclude that piecewise linear interpolations in time of the approximations $u^{m}_{h}$, $\widetilde{u}_{h}^{m}$ converge to the same limit due to the lack of the corresponding bounds in the energy inequality. Therefore, we provide an estimate on the time differences of $\widetilde{u}_{h}$ instead of $\partial_{t}u_{h}$, similar to [12]. This will be one of the ingredients to apply the Simon compactness result [26], restated in Theorem B.1, in order to derive compactness of the approximating sequences in $L^{2}([0,T]\times\Omega)$.
@@
Denote $Z=(H^{1}_{\operatorname{div}}(\Omega)\cap H^{s}(\Omega))^{*}$. Assume that $h^{s}\leq C\Delta t$, where $2\leq s\leq k+1$ and $k$ is the polynomial degree of $\mathbb{U}_{h}$. Then the aproximations $\widetilde{u}_{h}$ computed by (5.1)–(5.2) satisfy for any $0\leq\tau<T$,
@@
Given $\tau>0$ and $t\in(0,T-\tau]$, let $m_{1}\in\mathbb{N}$ such that $\widetilde{u}_{h}(t)=u_{h}^{m_{1}}$ and $m_{2}\in\mathbb{N}$ such that $\widetilde{u}_{h}(t+\tau)=\widetilde{u}_{h}^{m_{2}+1}$. If $\tau<\Delta t$, there is nothing to prove, so we can assume without loss of generality that $m_{2}\geq m_{1}$. Then
@@
We take a test function $v\in H^{1}_{\operatorname{div}}(\Omega)\cap H^{s}(\Omega)$, and plug in the scheme (5.1)–(5.2),
@@
$\leq\left\lVert\widetilde{u}^{m}_{h}\right\rVert_{L^{2}}\left\lVert\nabla\widetilde{u}_{h}^{m+1}\right\rVert_{L^{2}}\left\lVert\mathcal{P}_{\mathbb{U}_{h}}v\right\rVert_{L^{\infty}}+\frac{1}{2}\left\lVert\widetilde{u}^{m+1}_{h}\right\rVert_{L^{2}}\left\lVert\operatorname{div}\widetilde{u}_{h}^{m}\right\rVert_{L^{2}}\left\lVert\mathcal{P}_{\mathbb{U}_{h}}v\right\rVert_{L^{\infty}}$
$\leq C\left(\left\lVert\widetilde{u}^{m}_{h}\right\rVert_{L^{2}}\left\lVert\nabla\widetilde{u}_{h}^{m+1}\right\rVert_{L^{2}}+\left\lVert\widetilde{u}^{m+1}_{h}\right\rVert_{L^{2}}\left\lVert\nabla\widetilde{u}_{h}^{m}\right\rVert_{L^{2}}\right)\left\lVert v\right\rVert_{L^{\infty}}$
$\leq C\left(\left\lVert\widetilde{u}^{m}_{h}\right\rVert_{L^{2}}\left\lVert\nabla\widetilde{u}_{h}^{m+1}\right\rVert_{L^{2}}+\left\lVert\widetilde{u}^{m+1}_{h}\right\rVert_{L^{2}}\left\lVert\nabla\widetilde{u}_{h}^{m}\right\rVert_{L^{2}}\right)\left\lVert v\right\rVert_{H^{2}},$
@@
using (3.7b). For the third term, III, we have, using that $v$ is divergence free, the properties of the $L^{2}$-projection, (3.7), and that by the energy estimate $\left\lVert\nabla p\right\rVert_{L^{\infty}(0,T;L^{2}(\Omega))}\leq C\Delta t^{-1}$,
@@
$\leq C(\left\lVert\nabla p^{m}_{h}\right\rVert_{L^{2}(\Omega)}+\left\lVert\nabla p^{m-1}_{h}\right\rVert_{L^{2}(\Omega)})h^{s}\left\lVert v\right\rVert_{H^{s}(\Omega)}$
$\leq C\frac{h^{s}}{\Delta t}\left\lVert v\right\rVert_{H^{s}(\Omega)}$
$\leq C\left\lVert v\right\rVert_{H^{s}(\Omega)},$
under the condition that $h^{s}\leq C{\Delta t}$. Finally, for the last term, IV, we have
@@
Combining the estimates I – IV, we have, after taking the supremum over $v\in H^{1}_{\operatorname{div}}(\Omega)\cap H^{s}(\Omega)$,
$\left\lVert\widetilde{u}_{h}(t+\tau)-\widetilde{u}_{h}(t)\right\rVert_{Z}$ $\leq C\Delta t\sum_{m=m_{1}}^{m_{2}}\Big(\left\lVert\widetilde{u}^{m}_{h}\right\rVert_{L^{2}}\left\lVert\nabla\widetilde{u}_{h}^{m+1}\right\rVert_{L^{2}}+\left\lVert\widetilde{u}^{m+1}_{h}\right\rVert_{L^{2}}\left\lVert\nabla\widetilde{u}_{h}^{m}\right\rVert_{L^{2}}$
[…]
@@
$\left\lVert\widetilde{u}_{h}(t+\tau)-\widetilde{u}_{h}(t)\right\rVert_{Z}^{2}$ $\leq C(\tau+\Delta t)\Delta t\sum_{m=m_{1}}^{m_{2}}\Big(\left\lVert\widetilde{u}^{m}_{h}\right\rVert_{L^{2}}\left\lVert\nabla\widetilde{u}_{h}^{m+1}\right\rVert_{L^{2}}+\left\lVert\widetilde{u}^{m+1}_{h}\right\rVert_{L^{2}}\left\lVert\nabla\widetilde{u}_{h}^{m}\right\rVert_{L^{2}}$
@@
[…]
Next, we use Lemma 4.5 to show uniform time continuity of $\widetilde{u}_{h}$ in $L^{2}([0,T]\times\Omega)$ which we can then use together with Simon’s version of the Aubin-Lions-Simon lemma [26, Theorem 1], stated in Theorem B.1, to conclude precompactness of the sequence $\{\widetilde{u}_{h}\}_{h>0}$ in $L^{2}([0,T]\times\Omega)$. The following lemma and its proof closely follow [12, Lemma 3.6]. We include it here both for completeness and to illustrate that it can be established using Lemma 4.5, instead of the version employed in [12].